BALANCE SHEET ANALYSIS

Complete the balance sheet and sales information using the following financial data:

Total assets turnover: 1.1x
Days sales outstanding: 43 daysa
Inventory turnover ratio: 4x
Fixed assets turnover: 2.5x
Current ratio: 1.8x
Gross profit margin on sales: (Sales - Cost of goods sold)/Sales = 30%
aCalculation is based on a 365-day year. Do not round intermediate calculations. Round your answer to the nearest cent.

Balance Sheet
Cash $   Current liabilities $  
Accounts receivable    Long-term debt 80,000
Inventories    Common stock   
Fixed assets    Retained earnings 100,000
Total assets $400,000 Total liabilities and equity $  
Sales $   Cost of goods sold $  
Answer #1

Total Assets Turnover = Sales / Total Assets

1.1 = Sales / $400,000

Sales = 1.1 * $400,000 = $440,000

Gross profit margin on sales = (Sales - Cost of goods sold)/Sales

0.30 = ($440,000 - Cost of goods sold) / $440,000

0.30 * $440,000 = $440,000 - Cost of goods sold

Cost of goods sold = $440,000 - $132,000 = $308,000

Fixed Assets Turnover = Sales / Fixed Assets

2.5 = $440,000 / Fixed Assets

Fixed Assets = $440,000 / 2.5 = $176,000

Current Assets = Total Assets - Fixed Assets = $400,000 - $176,000 = $224,000

Current Ratio = Current Assets / Current Liabilities

1.8 = $224,000 / Current Liabilities

Current Liabilities = $224,000 / 1.8 = $124,444.44

Total Assets = Total Liabilities and equity = $400,000

Common Stock = Total Liabilities and equity - Current Liabilities - Long-term debt - Retained earnings

= $400,000 - $124,444.44 - $80,000 - $100,000 = $95,555.56

Inventory turnover ratio = Sales / Inventory

4 = $440,000 / Inventory

Inventory = $440,000 / 4 = $110,000

Days Sales Outstanding = 365 / [Sales / Accounts receivable]

43 = 365 / [$440,000 / Accounts Receivable]

$440,000 / Accounts Receivable = 365 / 43

Accounts Receivable = $440,000 / 8.49 = $51,835.62

Cash = Current Assets - Inventory - Accounts Receivable

= $224,000 - $110,000 - $51,835.62 = $62,164.38

Balance Sheet
Cash $62,164.38 Current liabilities $124,444.44
Accounts receivable $51,835.62 Long-term debt $80,000.00
Inventories $110,000.00 Common stock $95,555.56
Fixed assets $176,000.00 Retained earnings $100,000.00
Total assets $400,000.00 Total liabilities and equity $400,000.00
Sales $440,000.00 Cost of goods sold $308,000.00